### v1

# Gradient is All You Need: Gradient-Based Attention Fusion for Infrared Small Target Detection

Abstract
Infrared small target detection (IRSTD) is widely used in civilian and military applications. However, IRSTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds. To address this issue, we propose the Gradient Network (GaNet), which aims to extract and preserve edge and gradient information of small targets. GaNet employs the Gradient Transformer (GradFormer) module, simulating central difference convolutions (CDC) to extract and integrate gradient features with deeper features. Furthermore, we propose a global feature extraction model (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the background information. We compare the network with state-of-the-art (SOTA) approaches, and the results demonstrate that our method performs effectively. Our source code is available at https://github.com/greekinRoma/Gradient-Transformer.
Index Terms:
Infrared small target detection (IRSTD), convolution neural network (CNN), gradient Transformer, global feature extraction.
I Introduction
INFRARED small target detection (IRSTD) is critically important for numerous space-based computer vision applications, including remote sensing [1], precision guidance [2], and marine rescue [3]. While IRSTD has been extensively studied, it remains challenging. The considerable imaging distances associated with space-based infrared cameras often result in small targets that lack detailed texture and shape information.
Currently, the IRSTD methods can be categorized into model-driven and data-driven methods. Model-driven methods include three main approaches. 1) Filter-based methods, such as the Tophat [4] and Max-Median filters [5], enhance the target visibility in complex backgrounds while suppressing background noise. 2) Human visual system (HVS)-based methods, such as the local contrast measure (LCM)[6] and multi-patch contrast measure (MPCM)[7], leverage human visual perception to enhance local information. 3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch-image (IPI)[8] and reweight image patch tensor (RIPT)[9], decompose small targets and backgrounds using low-rank matrices. However, these techniques require manual parameter settings and often need help to adapt to complex scenarios, particularly in high noise levels or significant background clutter.
With the advancement of deep learning, data-driven approaches have achieved substantial progress in infrared small target detection (IRSTD). For example, the Asymmetric Context Modulation (ACM) network [10] introduces an asymmetric feature fusion technique, an alternative to the conventional skip connections in U-Net. Building on this, DNANet [11] implements a multi-layer nested architecture that supports progressive and adaptive interactions among feature layers. Moreover, UIUNet [12] enhances the detection of local target contrasts by integrating multiple U-Net structures and utilizing interactive cross-attention mechanisms for feature fusion. Additionally, GSTUnet [13] merges vision Transformer technology with CNNs in the encoder to efficiently extract edge features from small targets. Generally, these data-driven methods surpass traditional model-driven approaches. However, they have drawbacks, such as the substantial number of parameters and the potential loss of local feature information in higher layers, indicating opportunities for further enhancements.
In order to address the challenges of feature extraction, we propose the Gradient Attention Network (GaNet). GaNet incorporates two key innovation components to improve feature extraction. First, we propose a gradient Transformer (GradFormer) module that performs the feature gradient calculations by simulating the central difference convolutions (CDC) and enhances weak features to enable the network to learn a comprehensive feature representation of the target. Second, the global feature extraction module (GFEM) solves the problem of lacking global background perception, thereby improving the ability to obtain contextual information. The main contributions of this letter are as follows.
  1. 1.
    We propose the GANet, which utilizes the GradFormer to simulate the CDC with dynamic weights, extracting the edge and point information vital for IRSTD.
  2. 2.
    We propose the GFEM to improve the CNN from only focusing on local details to integrating these details with global information.
  3. 3.
    We conduct comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that GaNet outperforms existing methods.
II Methods
The overall architecture of GaNet is illustrated in Fig. 1. GradFormer is utilized to extract gradient information, and GFEM enhances contextual features. Finally, detailed information from low-level layers is integrated with high-level semantics with the global receptive field.
Figure 1: The overall structure of GaNet. The blue represents the downsampling module, the green indicates the GradFormer module, the orange corresponds to the upsampling module, the red signifies the concatenate operation, and the purple denotes the global feature extraction module (GFEM).
II-A Gradient Transformer (GradFormer)
CDCs are defined in Eq. 1 and are employed to extract edge and gradient features. Here, $c_{i}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$. While CDCs excel at extracting details, there is still room for improvement. A significant limitation is the use of fixed weights for different directions, meaning the importance of each direction remains constant across different images.
$out_{\hat{c}xy}=\sum^{c_{i}-1}_{i=0}\sum^{7}_{j=0}w_{\hat{c}ijxy}(b_{ijxy}-o_{ixy})\$ (1)
We can transform Eq. 1 into a fusion of Eq. 2 and Eq. 3. The process described in Eq. 2 is illustrated in Fig. 2. As shown in Fig. 2, we use edge convolution to extract the difference values from the input $T\in\mathbb{R}^{c_{i}\times w\times h}$ and reshape them to produce a local difference matrix $D\in\mathbb{R}^{8c_{i}\times wh}$. In Eq. 3, we utilize the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{i}}$ and the local difference matrix $D$ to obtain the final result $Out$. Here, $w$ denotes the width of the input, $h$ represents the input height, and $c_{o}$ signifies the number of output channels.
$D=reshape(Conv_{3\times 3}(T))\$ (2)
$Out=M_{w}D$ (3)
Figure 2: $D$ is obtained by performing edge detection convolution on $T_{i}$ and unfolding the output.
When $M_{w}$ is influenced by the input $T_{i}$, the CDC could exhibit enhanced robustness. Based on the analysis above, we introduce the GradFormer to simulate CDC and extract the details using dynamic weights from low-level layers for IRSTD. The structure of GradFormer is shown in Fig. 3.
Figure 3: The various heads in GradFormer employ the edge convolution to get $D_{\sum}$. Each head’s key and value are derived from $D_{\sum}$, while the query is generated from the flattened input.
In varied heads of the GradFormer, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
In the various heads of GradFormer, we apply different edge detection convolutions from Fig. 2 to $T$, utilizing varying dilation ratios. The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{i}\times wh}$. Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{i}\times wh}$. The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{i}\times wh}$ functions as both the key and the value.
$Q^{n}$ $=W^{n}_{Q}O,K^{n}=W^{n}_{K}D^{n},V=W^{n}_{V}D^{n}$ (4)
where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{i}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{i}}{m}\times 8c_{i}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{i}}{m}\times 8c_{i}}$ are the reshaped weights of convolutions, $n$ denotes dilation ratios, and $m$ denotes the number of heads. We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{i}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{i}}{m}}$. $V^{n}\in\mathbb{R}^{\frac{8c_{i}}{m}\times wh}$ denotes the value of the transformer. The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:
$Out^{n}$ $=Softmax(Norm(\frac{Q^{n}(K^{n})^{T}}{w}))V^{n}$ (5)
$=M^{n}V^{n}=(M^{n}W_{V}^{n})D^{n}=M^{n}_{mix}D^{n}$
where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{i}}$ denotes the dynamic matrix derived from a attention matrix $M^{n}$ and a fixed convolution weight matrix $W_{V}^{n}$. According to Eq. 3, $Out^{n}$ is the equivalent result of the CDC with dynamic weights and a dilation ratio of n. $Norm$ is the Instance Normalization. At last, we concatenate the results in varied heads, reshape them and use $1\times 1$ convolution to fuse them as shown in the Equation below:
$Out=Conv_{1\times 1}(reshape(Out^{1},...,Out^{m}))$ (6)
Overall, the gradient information from features at different scales can be fully extracted, enhancing the model’s ability to adapt to complex backgrounds.
II-B Global Feature Extraction Module
Background information is crucial alongside detailed features in IRSTD. However, perceiving background relies on global features, which the GradFormer struggles to handle effectively. To address this limitation and enhance global information perception, we propose the Global Feature Enhancement Module (GFEM), as illustrated in Fig. 4.
In GFEM, we utilize the attention mechanism to facilitate the transmission of information across distant pixels. This module includes two key aspects to broaden perception: First, we employ the non-local attention mechanism [14] to compute the spatial attention, enabling connectivity between pixels over long distances. Second, we incorporate a squeeze-and-excitation block [15] to enhance channel-wise perception. Finally, we concatenate the outputs from the global spatial and channel attention modules and apply convolution to fuse them, enabling GFEM to capture comprehensive feature representations.
Subsequently, each feature map acquires both spatial and channel global receptive fields, which are vital for effective object detection.
Figure 4: Diagram of the Global Feature Extraction Module (GFEM) using the Non-local Attention Module to capture global spatial features and Squeeze-and-Excitation Block to extract global channel information.
II-C Loss Function
The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, where the predictions of GaNet are mapped into the range of 0 and 1, using the Sigmoid function as shown in Eq. 7.
$p_{i,j}=\frac{1}{1+e^{-v_{i,j}}}$ (7)
where $p_{i,j}$ denotes the Sigmoid output of $v_{i,j}$, the model’s prediction at pixel $(i,j)$. Moreover, the SoftIoU loss is defined as:
$SoftIoU=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (8)
$SoftIoU_{loss}=1-SoftIoU$ (9)
In this equation, $g_{i,j}$ denotes the ground truth value at pixel $(i,j)$, typically 0 or 1.
Figure 5: Visual results from varied data-driven methods. Red, green, and yellow boxes label detected targets, missed targets, and false alarms.
III Experiment and Analysis
This section describes the experimental details, including a series of ablation studies to verify the proposed modules. Comparative experiments with other methods demonstrate that our approach outperforms existing state-of-the-art (SOTA) methods.
III-A Dataset and Evaluation Metrics
The datasets used for evaluating the GaNet are NUDT-SIRST[11] and IRSTD-1k[16], which are widely employed to assess model performance. The evaluation metrics include mean intersection over union ($mIoU$), F-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$).
III-B Experimental Environment and Parameter Settings
Our model is implemented by the PyTorch framework on NVIDIA GeForce RTX 4080 GPUs. We employ the Adam optimizer, starting with an initial learning rate of $5\times 10^{-4}$. This learning rate is reduced to $5\times 10^{-5}$ at epoch 200 and further decrease to $5\times 10^{-6}$ We use a batch size of 4 for training and run the model for 400 epochs.
III-C Ablation Study
TABLE I: Studies on different dilation rates on NUDT-SIRST (The best results are bold)
Dilation Rate mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
1 93.53 96.30 98.89 5.58
1,3 94.82 97.34 98.94 1.82
1,5 94.24 97.03 98.65 1.47
TABLE II: Studies Of Different Components on NUDT-SIRST (The best results are bold)
Module mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
U-Net 91.31 95.44 97.98 4.46
U-Net+GradFormer 94.25 97.03 98.83 2.73
U-Net+GFEM 92.32 96.14 96.30 3.96
U-Net+GradFormer+GFEM 94.82 97.34 98.94 1.82
III-C1 Studies of Dilation Rate for GradFormer
As mentioned in II-A, GradFormer employs varying dilation rates across different heads. Tab. I demonstrates the efficacy of this strategy, as models featuring diverse dilation ratios outperform those with a uniform rate. Furthermore, networks employing dilation ratios of 1 and 3 yield superior results compared to those with larger ratios, indicating that smaller dilation ratios are more effective for capturing local differences in infrared small targets, likely due to their small size. Therefore, the dilation ratio significantly impacts target detection, making its configuration crucial for the effectiveness of GradFormer.
TABLE III: Quantitative Comparsion With Different Methods (The best results are bold, second best results are underline.)
Module Metrics NUDT-SIRST IRSTD1K
Params(M) mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$ mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$
ACM[10] 0.40 69.69 82.10 96.82 8.12 63.78 77.91 93.93 5.81
RDIAN[17] 0.90 86.67 93.81 97.98 4.76 64.08 78.12 93.34 5.30
ALCNet[18] 0.37 92.39 96.04 98.83 3.79 65.10 78.88 93.26 7.90
AGPCNet[19] 13.36 86.71 92.88 98.09 3.93 67.54 80.64 94.27 4.76
DNANet[11] 4.69 93.41 96.59 98.94 3.52 65.79 79.23 93.60 5.96
UIU-Net[12] 50.54 94.27 97.04 98.73 1.22 68.06 80.91 92.59 5.85
SCTransNet[20] 11.19 93.38 96.57 97.88 4.73 67.97 80.90 93.26 6.24
Ours 4.46 94.82 97.34 98.94 1.82 68.26 81.11 95.95 4.42
III-C2 Studies of Various Components
We illustrate the effectiveness of each module in Tab. II. The performance of our network improves progressively with the inclusion of each module. GradFormer, which focuses on local difference features, enhances the model’s performance more effectively than GFEM, which primarily relies on global perception. In IRSTD, the capability to learn gradient information is more critical than access to global context.
III-D Comparsion with State-of-the-art (SOTA) Methods
To demonstrate the effectiveness of our proposed model, we compare it with several SOTA methods on two datasets, including SCTransNet [20], UIUNet [12], ACM [10], ALCNet [18], AGPCNet [19], RDIAN [17], and DNANet [11]. As shown in Tab. III, and our method performs excellently. We utilize GaNet with dilation ratios of 1 and 5 as the baseline models for comparative evaluation.
III-D1 Qualitative Result
As shown in Fig. 5, our model preserves the shape of targets with low missed detections (green boxes) and false alarms (yellow boxes) compared to other methods. Our approach excels in learning rich feature representations by seamlessly integrating detailed features with high-level semantics and leveraging a global attention, which enables precise differentiation between targets and backgrounds, effectively mitigating false alarms and reducing missed detections. Despite the challenges posed by small and dim targets in infrared images with complex backgrounds, our network consistently balances segmentation quality with detection accuracy. While methods like UIU-Net may retain target structure, they often suffer from false positives and negatives. Other approaches, such as ALCNet, may exhibit fewer false alarms and missed targets but struggle to maintain target shapes. Through comparisons with other methods, we have highlighted the superior performance of our approach.
III-D2 Quantitative Results
As shown in Tab. III, our method achieves impressive performance, including an $mIoU$ of $95.21\%$, an $F1$ score of $97.51\%$, and a $Pd$ of $99.04\%$, a $Pa$ of $1.06\times 10^{-6}$. With $4.46$M parameters, it is significantly more efficient than other methods. Similarly, on the IRSTD-1k dataset, our method is a leading network. Generally, methods that rely on data with more complex structures tend to perform better. However, our network demonstrates that such models often extract features that are irrelevant to IRSTD. Instead, we focus on utilizing gradients from lower layers and semantics from deeper layers for optimal performance. Additionally, we employ the ROC curve to analyze the performance of various models, as depicted in Fig. 6. While UIU-Net achieves commendable results in $mIoU$, its AUC scores are comparatively weak. In contrast, ALCNet’s performance shows significant divergence, with subpar outcomes in $mIoU$ but relatively strong performance in ROC metrics. The model proposed in this article consistently demonstrates good performance in both ROC and $mIoU$.
Figure 6: ROC curves of varied networks on NUDT-SIRST.
IV Conclusion
In this letter, we propose an IRSTD network that enhances detection performance compared to baseline models and has fewer model parameters. The GradFormer module was proposed to improve the detection performance of the network. Additionally, we propose the GFEM module, which focuses on global perception across the entire feature map and learns relationships between distant pixels, preventing GradFormer from overlooking background information due to its emphasis on local differences. To demonstrate the effectiveness of the proposed network and modules, we conducted extensive experiments that indicate satisfied results and significant practical application potential.

### v2

@@
-Gradient is All You Need: Gradient-Based Attention Fusion for Infrared Small Target Detection
+DATransNet: Dynamic Attention Transformer Network for Infrared Small Target Detection
 Abstract
-Infrared small target detection (IRSTD) is widely used in civilian and military applications.
-However, IRSTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds.
-To address this issue, we propose the Gradient Network (GaNet), which aims to extract and preserve edge and gradient information of small targets.
-GaNet employs the Gradient Transformer (GradFormer) module, simulating central difference convolutions (CDC) to extract and integrate gradient features with deeper features.
-Furthermore, we propose a global feature extraction model (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the background information.
+Infrared small target detection (ISTD) is widely used in civilian and military applications.
+However, ISTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds.
+To address this issue, we propose the Dynamic Attention Transformer Network (DATransNet), which aims to extract and preserve edge information of small targets.
+DATransNet employs the Dynamic Attention Transformer (DATrans), simulating central difference convolutions (CDC) to extract and integrate gradient features with deeper features.
+Furthermore, we propose a global feature extraction module (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the background information.
 We compare the network with state-of-the-art (SOTA) approaches, and the results demonstrate that our method performs effectively.
-Our source code is available at https://github.com/greekinRoma/Gradient-Transformer.
+Our source code is available at https://github.com/greekinRoma/DATransNet.
 Index Terms:
-Infrared small target detection (IRSTD), convolution neural network (CNN), gradient Transformer, global feature extraction.
+Infrared small target detection (ISTD), convolution neural network (CNN), Dynamic Attention Transformer, global feature extraction.
 I Introduction
-INFRARED small target detection (IRSTD) is critically important for numerous space-based computer vision applications, including remote sensing [1], precision guidance [2], and marine rescue [3].
-While IRSTD has been extensively studied, it remains challenging.
+INFRARED small target detection (ISTD) is critically important for numerous space-based computer vision applications.
+Although ISTD has been extensively studied, it remains a challenge.
 The considerable imaging distances associated with space-based infrared cameras often result in small targets that lack detailed texture and shape information.
-Currently, the IRSTD methods can be categorized into model-driven and data-driven methods.
+Currently, the ISTD methods can be categorized into model-driven and data-driven methods.
 Model-driven methods include three main approaches.
-1) Filter-based methods, such as the Tophat [4] and Max-Median filters [5], enhance the target visibility in complex backgrounds while suppressing background noise.
-2) Human visual system (HVS)-based methods, such as the local contrast measure (LCM)[6] and multi-patch contrast measure (MPCM)[7], leverage human visual perception to enhance local information.
-3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch-image (IPI)[8] and reweight image patch tensor (RIPT)[9], decompose small targets and backgrounds using low-rank matrices.
-However, these techniques require manual parameter settings and often need help to adapt to complex scenarios, particularly in high noise levels or significant background clutter.
-With the advancement of deep learning, data-driven approaches have achieved substantial progress in infrared small target detection (IRSTD).
-For example, the Asymmetric Context Modulation (ACM) network [10] introduces an asymmetric feature fusion technique, an alternative to the conventional skip connections in U-Net.
-Building on this, DNANet [11] implements a multi-layer nested architecture that supports progressive and adaptive interactions among feature layers.
-Moreover, UIUNet [12] enhances the detection of local target contrasts by integrating multiple U-Net structures and utilizing interactive cross-attention mechanisms for feature fusion.
+1) Filter-based methods, such as the Tophat [1], enhance target visibility in complex backgrounds while suppressing background noise.
+2) Methods based on the human visual system (HVS), such as the local contrast measure (LCM)[2] and multi-patch contrast measure (MPCM)[3], leverage human visual perception to improve local information.
+3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch image (IPI)[4], and Rewight infrared patch tensor (RIPT) [5], decompose small targets and backgrounds using low-rank matrices.
+However, these techniques require manual parameter settings and often struggle to adapt to complex scenarios.
+With the advancement of deep learning, data-driven approaches have achieved substantial progress in infrared small target detection (ISTD).
+For example, the Asymmetric Context Modulation (ACM) network [6] introduces an asymmetric feature fusion technique, an alternative to conventional skip connections in U-Net.
+Building on this, DNANet [7] implements a multilayer nested architecture that supports progressive and adaptive interactions among feature layers.
+Moreover, UIUNet [8] enhances the detection of local target contrasts by integrating multiple U-Net structures and using interactive cross-attention mechanisms for feature fusion.
 Additionally, GSTUnet [9] merges vision Transformer technology with CNNs in the encoder to efficiently extract edge features from small targets.
 Generally, these data-driven methods surpass traditional model-driven approaches.
-However, they have drawbacks, such as the substantial number of parameters and the potential loss of local feature information in higher layers, indicating opportunities for further enhancements.
-In order to address the challenges of feature extraction, we propose the Gradient Attention Network (GaNet).
-GaNet incorporates two key innovation components to improve feature extraction.
-First, we propose a gradient Transformer (GradFormer) module that performs the feature gradient calculations by simulating the central difference convolutions (CDC) and enhances weak features to enable the network to learn a comprehensive feature representation of the target.
+Some challenges faced by data-driven methods, such as the substantial number of parameters and the potential loss of local feature information.
+Our motivation focus on enhancing the feature information from small targets.
+Although Data-driven methods achieve good performance, they suffer from a substantial number of parameters and potential loss of local information.
+Therefore, we propose some methods to improve the ISTD network, and the new network is called Dynamic Attention Transformer Network (DATransNet).
+DATransNet incorporates two key innovation components to improve feature extraction.
+First, we propose the Dynamic Attention Transformer (DATrans), which utilizes Transformers to simulate the central difference convolution (CDC) and enable the network to retain the detailed information of the inputs in deep layers and give dynamic weights to differences result from differenct directions.
 Second, the global feature extraction module (GFEM) solves the problem of lacking global background perception, thereby improving the ability to obtain contextual information.
@@
   1. 1.
-    We propose the GANet, which utilizes the GradFormer to simulate the CDC with dynamic weights, extracting the edge and point information vital for IRSTD.
+    We propose the DATransNet, which utilizes the DATrans to simulate the CDC with dynamic weights, extracting the edge and point information.
   2. 2.
@@
   3. 3.
-    We conduct comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that GaNet outperforms existing methods.
+    We conduct comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that DATransNet outperforms existing methods.
 II Methods
-The overall architecture of GaNet is illustrated in Fig.
-1.
-GradFormer is utilized to extract gradient information, and GFEM enhances contextual features.
-Finally, detailed information from low-level layers is integrated with high-level semantics with the global receptive field.
-Figure 1: The overall structure of GaNet.
-The blue represents the downsampling module, the green indicates the GradFormer module, the orange corresponds to the upsampling module, the red signifies the concatenate operation, and the purple denotes the global feature extraction module (GFEM).
-II-A Gradient Transformer (GradFormer)
+The overall architecture of DATransNet is illustrated in Fig. 1. DATrans is based on U-Net and implements DATrans and GFEM to extract detailed information and contextual features.
+Figure 1: The overall structure of DATransNet. The red represents the downsampling module and the blue corresponds to the upsampling module.
+II-A Dynamic Attention Transformer (DATrans)
 CDCs are defined in Eq.
 1 and are employed to extract edge and gradient features.
-Here, $c_{i}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$.
-While CDCs excel at extracting details, there is still room for improvement.
-A significant limitation is the use of fixed weights for different directions, meaning the importance of each direction remains constant across different images.
-$out_{\hat{c}xy}=\sum^{c_{i}-1}_{i=0}\sum^{7}_{j=0}w_{\hat{c}ijxy}(b_{ijxy}-o_{ixy})\$ (1)
+Here, $c_{inp}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$.
+$out_{\hat{c}xy}$ denotes the output of the $\hat{c}$-th channel at position $(x,y)$.
+$w_{\hat{c}ijxy}$ is the weight applied to the difference $(o_{ixy}-b_{ijxy})$, which contributes to determining $out_{\hat{c}xy}$.
+A significant limitation of CDCs is the use of fixed weights for different directions, meaning the importance of each direction remains constant across different images.
+$out_{\hat{c}xy}=\sum^{c_{inp}-1}_{i=0}\sum^{7}_{j=0}w_{\hat{c}ijxy}(b_{ijxy}-o_{ixy})\$ (1)
 We can transform Eq.
@@
 As shown in Fig.
-2, we use edge convolution to extract the difference values from the input $T\in\mathbb{R}^{c_{i}\times w\times h}$ and reshape them to produce a local difference matrix $D\in\mathbb{R}^{8c_{i}\times wh}$.
+2, we use edge convolution to extract the difference values from the input $T\in\mathbb{R}^{c_{inp}\times w\times h}$ and reshape them to produce a local difference matrix $D\in\mathbb{R}^{8c_{inp}\times wh}$.
 In Eq.
-3, we utilize the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{i}}$ and the local difference matrix $D$ to obtain the final result $Out$.
+3, we utilize the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{inp}}$ and the local difference matrix $D$ to obtain the final result $Out$.
 Here, $w$ denotes the width of the input, $h$ represents the input height, and $c_{o}$ signifies the number of output channels.
@@
 $Out=M_{w}D$ (3)
-Figure 2: $D$ is obtained by performing edge detection convolution on $T_{i}$ and unfolding the output.
-When $M_{w}$ is influenced by the input $T_{i}$, the CDC could exhibit enhanced robustness.
-Based on the analysis above, we introduce the GradFormer to simulate CDC and extract the details using dynamic weights from low-level layers for IRSTD.
-The structure of GradFormer is shown in Fig.
+Figure 2: $D$ is obtained by performing edge detection convolution on $T$ and unfolding the output.
+When $M_{w}$ is influenced by the input $T$, the CDC could exhibit enhanced robustness.
+Based on the analysis above, we introduce the DATrans to simulate CDC and extract the details using dynamic weights from low-level layers for ISTD.
+The structure of DATrans is shown in Fig.
 3.
-Figure 3: The various heads in GradFormer employ the edge convolution to get $D_{\sum}$. Each head’s key and value are derived from $D_{\sum}$, while the query is generated from the flattened input.
-In varied heads of the GradFormer, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
-In the various heads of GradFormer, we apply different edge detection convolutions from Fig.
+Figure 3: The various heads in DATrans employ the edge convolution to get $D_{\sum}$. Each head’s key and value are derived from $D_{\sum}$, while the query is generated from the flattened input.
+In varied heads of the DATrans, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
+In the various heads of DATrans, we apply different edge detection convolutions from Fig.
 2 to $T$, utilizing varying dilation ratios.
-The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{i}\times wh}$.
-Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{i}\times wh}$.
-The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{i}\times wh}$ functions as both the key and the value.
+The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$.
+Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$.
+The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ functions as both the key and the value.
 $Q^{n}$ $=W^{n}_{Q}O,K^{n}=W^{n}_{K}D^{n},V=W^{n}_{V}D^{n}$ (4)
-where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{i}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{i}}{m}\times 8c_{i}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{i}}{m}\times 8c_{i}}$ are the reshaped weights of convolutions, $n$ denotes dilation ratios, and $m$ denotes the number of heads.
-We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{i}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{i}}{m}}$.
-$V^{n}\in\mathbb{R}^{\frac{8c_{i}}{m}\times wh}$ denotes the value of the transformer.
+where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions, $n$ denotes dilation ratios, and $m$ denotes the number of heads.
+We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$.
+$V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ denotes the value of the transformer.
 The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:
@@
 $=M^{n}V^{n}=(M^{n}W_{V}^{n})D^{n}=M^{n}_{mix}D^{n}$
-where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{i}}$ denotes the dynamic matrix derived from a attention matrix $M^{n}$ and a fixed convolution weight matrix $W_{V}^{n}$.
+where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{inp}}$ denotes the dynamic matrix derived from a attention matrix $M^{n}$ and a fixed convolution weight matrix $W_{V}^{n}$.
 According to Eq.
@@
 Overall, the gradient information from features at different scales can be fully extracted, enhancing the model’s ability to adapt to complex backgrounds.
-II-B Global Feature Extraction Module
-Background information is crucial alongside detailed features in IRSTD.
-However, perceiving background relies on global features, which the GradFormer struggles to handle effectively.
-To address this limitation and enhance global information perception, we propose the Global Feature Enhancement Module (GFEM), as illustrated in Fig.
+II-B Global Feature Extraction Module (GFEM)
+Background information is crucial alongside detailed features in ISTD.
+However, perceiving background relies on global features, which the DATrans struggles to handle effectively.
+To address this limitation and enhance global information perception, we propose the GFEM, as illustrated in Fig.
 4.
@@
 Subsequently, each feature map acquires both spatial and channel global receptive fields, which are vital for effective object detection.
-Figure 4: Diagram of the Global Feature Extraction Module (GFEM) using the Non-local Attention Module to capture global spatial features and Squeeze-and-Excitation Block to extract global channel information.
+Figure 4: The structure of the GFEM using the Non-local (Non-local Attention Module) to capture global spatial features and SE Block (Squeeze-and-Excitation Block) to extract global channel information.
 II-C Loss Function
-The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, where the predictions of GaNet are mapped into the range of 0 and 1, using the Sigmoid function as shown in Eq.
+The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, where the predictions of DATransNet are mapped into the range of 0 and 1, using the Sigmoid function as shown in Eq.
 7.
@@
 where $p_{i,j}$ denotes the Sigmoid output of $v_{i,j}$, the model’s prediction at pixel $(i,j)$. Moreover, the SoftIoU loss is defined as:
-$SoftIoU=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (8)
-$SoftIoU_{loss}=1-SoftIoU$ (9)
+$loss=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (8)
 In this equation, $g_{i,j}$ denotes the ground truth value at pixel $(i,j)$, typically 0 or 1.
-Figure 5: Visual results from varied data-driven methods. Red, green, and yellow boxes label detected targets, missed targets, and false alarms.
+Figure 5: Visual results from varied data-driven methods.
+Red, green, and yellow boxes label detected targets, missed targets, and false alarms.
+The Red bounding box are zoom in, facilitating to auxiliary our to evaluation detection results.
+Our DATransNet can achieve high probability detection rate, and achieve precision localization and segmentation of their shapes.
 III Experiment and Analysis
-This section describes the experimental details, including a series of ablation studies to verify the proposed modules.
-Comparative experiments with other methods demonstrate that our approach outperforms existing state-of-the-art (SOTA) methods.
+This section first describes the experimental details and the evaluation metrics.
+Following with a series of ablation studies to verify the proposed modules.
+Finally, a comprehensive comparative experiments with other methods on qualitative and quantitative results demonstrate that our approach outperforms existing state-of-the-art (SOTA) methods.
 III-A Dataset and Evaluation Metrics
-The datasets used for evaluating the GaNet are NUDT-SIRST[11] and IRSTD-1k[16], which are widely employed to assess model performance.
+The datasets used for evaluating the DATransNet are NUDT-SIRST[7] and IRSTD-1K[12].
 The evaluation metrics include mean intersection over union ($mIoU$), F-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$).
+Meanwhile, we plot the ROC curves of various methods to evaluate the robust of DATransNet.
 III-B Experimental Environment and Parameter Settings
-Our model is implemented by the PyTorch framework on NVIDIA GeForce RTX 4080 GPUs.
+Our model is implemented by PyTorch framework, then training on NVIDIA GeForce RTX 4080 GPUs.
 We employ the Adam optimizer, starting with an initial learning rate of $5\times 10^{-4}$.
@@
 III-C Ablation Study
-TABLE I: Studies on different dilation rates on NUDT-SIRST (The best results are bold)
+III-C1 Module-wise Performance Gain
+In this ablation study, we begin with the most primitive Baseline, Which includes only feature extraction and upsampling with in the U-Net framework.
+Then, We test the performance of the DATransNet w/o each module, the results are shown in Tab.
+I.
+The results illustrate the effectiveness of each module.
+The performance of our network improves progressively with the inclusion of DATrans and GFEM.
+DATrans, which focuses on local difference features, enhances the model’s performance more effectively than GFEM, which primarily relies on global perception.
+In ISTD, the capability to learn gradient information is more critical than access to global context.
+TABLE I: Studies Of Different Components on NUDT-SIRST (The best results are bold)
+Module mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
+U-Net 91.31 95.44 97.98 4.46
+U-Net+DATrans 94.25 97.03 98.83 2.73
+U-Net+GFEM 92.32 96.14 96.30 3.96
+U-Net+DATrans+GFEM 94.93 97.39 99.04 2.00
+III-C2 Studies of Dilation Rate for DATrans
+As mentioned in Section II-A, DATrans employs varying dilation rates across different detection heads.
+Tab.
+II demonstrates the efficacy of this strategy, as models with diverse dilation ratios outperform those with a uniform rate.
+Furthermore, networks with dilation ratios of 1 and 3 yield superior results compared to those with other dilation ratios.
+Therefore, the dilation ratio significantly impacts target detection, and we choose the ratios of 1 and 3 as the dilation ratios for our network.
+TABLE II: Studies on different dilation rates on NUDT-SIRST (The best results are bold)
 Dilation Rate mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
 1 93.53 96.30 98.89 5.58
-1,3 94.82 97.34 98.94 1.82
+1,2 94.41 97.12 98.83 2.46
+1,3 94.93 97.39 99.04 2.00
 1,5 94.24 97.03 98.65 1.47
-TABLE II: Studies Of Different Components on NUDT-SIRST (The best results are bold)
-Module mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
-U-Net 91.31 95.44 97.98 4.46
-U-Net+GradFormer 94.25 97.03 98.83 2.73
-U-Net+GFEM 92.32 96.14 96.30 3.96
-U-Net+GradFormer+GFEM 94.82 97.34 98.94 1.82
-III-C1 Studies of Dilation Rate for GradFormer
-As mentioned in II-A, GradFormer employs varying dilation rates across different heads.
-Tab.
-I demonstrates the efficacy of this strategy, as models featuring diverse dilation ratios outperform those with a uniform rate.
-Furthermore, networks employing dilation ratios of 1 and 3 yield superior results compared to those with larger ratios, indicating that smaller dilation ratios are more effective for capturing local differences in infrared small targets, likely due to their small size.
-Therefore, the dilation ratio significantly impacts target detection, making its configuration crucial for the effectiveness of GradFormer.
-TABLE III: Quantitative Comparsion With Different Methods (The best results are bold, second best results are underline.)
-Module Metrics NUDT-SIRST IRSTD1K
-Params(M) mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$ mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$
-ACM[10] 0.40 69.69 82.10 96.82 8.12 63.78 77.91 93.93 5.81
-RDIAN[17] 0.90 86.67 93.81 97.98 4.76 64.08 78.12 93.34 5.30
-ALCNet[18] 0.37 92.39 96.04 98.83 3.79 65.10 78.88 93.26 7.90
-AGPCNet[19] 13.36 86.71 92.88 98.09 3.93 67.54 80.64 94.27 4.76
-DNANet[11] 4.69 93.41 96.59 98.94 3.52 65.79 79.23 93.60 5.96
-UIU-Net[12] 50.54 94.27 97.04 98.73 1.22 68.06 80.91 92.59 5.85
-SCTransNet[20] 11.19 93.38 96.57 97.88 4.73 67.97 80.90 93.26 6.24
-Ours 4.46 94.82 97.34 98.94 1.82 68.26 81.11 95.95 4.42
-III-C2 Studies of Various Components
-We illustrate the effectiveness of each module in Tab.
-II.
-The performance of our network improves progressively with the inclusion of each module.
-GradFormer, which focuses on local difference features, enhances the model’s performance more effectively than GFEM, which primarily relies on global perception.
-In IRSTD, the capability to learn gradient information is more critical than access to global context.
+1,2,3,4 94.58 97.20 98.04 2.21
+TABLE III: Studies on GFEM on NUDT-SIRST (The best results are bold)
+Non-local SE Block mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
+$\checkmark$ -- 94.69 97.27 98.73 2.57
+-- $\checkmark$ 94.53 97.19 98.23 1.75
+$\checkmark$ $\checkmark$ 94.93 97.39 99.04 2.00
+III-C3 Studies of GFEM
+As mentioned in III, we conduct the experiments on the modules in the GFEM. As we can see, the network combining the two modules performs better.
 III-D Comparsion with State-of-the-art (SOTA) Methods
-To demonstrate the effectiveness of our proposed model, we compare it with several SOTA methods on two datasets, including SCTransNet [20], UIUNet [12], ACM [10], ALCNet [18], AGPCNet [19], RDIAN [17], and DNANet [11].
+To evaluate the effectiveness of our proposed model, we compare it with several SOTA methods on NUDT-SIRST and IRSTD-1K datasets, including SCTransNet [16], UIUNet [8], ACM [6], ALCNet [14], AGPCNet [15], RDIAN [13], and DNANet [7].
 As shown in Tab.
-III, and our method performs excellently.
-We utilize GaNet with dilation ratios of 1 and 5 as the baseline models for comparative evaluation.
+IV, and our method performs excellently.
+We utilize DATransNet with dilation ratios of 1 and 5 as the baseline models for comparative evaluation.
+The comparison experiments include qualitative results and quantitative results, receptively.
 III-D1 Qualitative Result
 As shown in Fig.
-5, our model preserves the shape of targets with low missed detections (green boxes) and false alarms (yellow boxes) compared to other methods.
-Our approach excels in learning rich feature representations by seamlessly integrating detailed features with high-level semantics and leveraging a global attention, which enables precise differentiation between targets and backgrounds, effectively mitigating false alarms and reducing missed detections.
-Despite the challenges posed by small and dim targets in infrared images with complex backgrounds, our network consistently balances segmentation quality with detection accuracy.
-While methods like UIU-Net may retain target structure, they often suffer from false positives and negatives.
-Other approaches, such as ALCNet, may exhibit fewer false alarms and missed targets but struggle to maintain target shapes.
-Through comparisons with other methods, we have highlighted the superior performance of our approach.
+5, our model preserves the shape of targets more closely to ground truth with low missed detections (green boxes) and false alarms (yellow boxes) compared to other methods, achieve satisfied results.
+TABLE IV: Quantitative Comparsion With Different Methods on NUDT-SIRST and IRSTD-1K (The best results are bold, second best results are underline.)
+Module Metrics NUDT-SIRST IRSTD-1K
+Params(M) mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$ mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$
+ACM[6] 0.40 70.97 82.99 97.67 7.33 63.78 77.91 93.93 5.81
+RDIAN[13] 0.90 87.78 93.47 97.88 9.67 64.08 78.12 93.34 5.30
+ALCNet[14] 0.37 92.45 96.08 98.94 2.62 65.10 78.88 93.26 7.90
+AGPCNet[15] 13.36 85.38 92.11 97.88 5.95 67.54 80.64 94.27 4.76
+DNANet[7] 4.69 93.73 96.70 99.25 4.55 65.79 79.23 93.60 5.96
+UIU-Net[8] 50.54 94.11 96.96 97.98 0.74 68.06 80.91 92.59 5.85
+SCTransNet[16] 11.19 93.83 96.77 98.20 1.56 67.97 80.90 93.26 6.24
+Ours 4.46 94.93 97.39 99.04 2.00 68.26 81.11 95.95 4.42
 III-D2 Quantitative Results
-As shown in Tab.
-III, our method achieves impressive performance, including an $mIoU$ of $95.21\%$, an $F1$ score of $97.51\%$, and a $Pd$ of $99.04\%$, a $Pa$ of $1.06\times 10^{-6}$.
-With $4.46$M parameters, it is significantly more efficient than other methods.
+As shown in the Tab.
+IV, our method achieves impressive performance, including an $mIoU$ of $94.93\%$, an $F1$ score of $97.39\%$, and a $Pd$ of $99.04\%$, a $Pa$ of $2.00\times 10^{-6}$.
 Similarly, on the IRSTD-1K dataset, our method is a leading network.
-Generally, methods that rely on data with more complex structures tend to perform better.
-However, our network demonstrates that such models often extract features that are irrelevant to IRSTD.
-Instead, we focus on utilizing gradients from lower layers and semantics from deeper layers for optimal performance.
-Additionally, we employ the ROC curve to analyze the performance of various models, as depicted in Fig.
+Figure 6: ROC curves of varied networks on NUDT-SIRST datasets.
+Furthermore, we used the ROC curve to analyze the performance of various models, as depicted in Fig.
 6.
-While UIU-Net achieves commendable results in $mIoU$, its AUC scores are comparatively weak.
-In contrast, ALCNet’s performance shows significant divergence, with subpar outcomes in $mIoU$ but relatively strong performance in ROC metrics.
-The model proposed in this article consistently demonstrates good performance in both ROC and $mIoU$.
-Figure 6: ROC curves of varied networks on NUDT-SIRST.
+Although UIU-Net achieves commendable results in $mIoU$, its AUC scores are comparatively weak.
+In contrast, ALCNet performance shows a significant divergence, with subpar outcomes in $mIoU$ but relatively strong performance in ROC metrics.
+The ROC curve corroborates the findings in Tab.
+IV.
+The model proposed in this article consistently shows good performance in both ROC and $mIoU$.
 IV Conclusion
-In this letter, we propose an IRSTD network that enhances detection performance compared to baseline models and has fewer model parameters.
-The GradFormer module was proposed to improve the detection performance of the network.
-Additionally, we propose the GFEM module, which focuses on global perception across the entire feature map and learns relationships between distant pixels, preventing GradFormer from overlooking background information due to its emphasis on local differences.
-To demonstrate the effectiveness of the proposed network and modules, we conducted extensive experiments that indicate satisfied results and significant practical application potential.
+In this letter, we propose an ISTD network that enhances detection performance.
+The DATrans was proposed to enhance the local gradient feature extraction and detection performance of the network.
+Additionally, we propose the GFEM, which focuses on global perception across the entire feature map and learns relationships between distant pixels, preventing DATrans from overlooking background information due to its emphasis on local gradient differences.
+Then the local and global feature are fused.
+We conducted extensive experiments indicating satisfied results with fewer parameters.
+The model is a back-box approach because its capability for dynamic weight adjustments lacks mathematical proof.
+Moreover, it suffers from false alarms when processing high-noise images.
+To address these limitations, we aim to improve the model in two aspects.
+At first, we plan to improve the interpretability of the network.
+Then, we would like to integrate temporal information into the model to improve its robustness.

### v3

@@
 The main contributions of this letter are as follows.
+Figure 1: The overall structure of DATransNet. The red represents the downsampling module and the blue corresponds to the upsampling module.
   1. 1.
@@
 The overall architecture of DATransNet is illustrated in Fig. 1. DATrans is based on U-Net and implements DATrans and GFEM to extract detailed information and contextual features.
-Figure 1: The overall structure of DATransNet. The red represents the downsampling module and the blue corresponds to the upsampling module.
 II-A Dynamic Attention Transformer (DATrans)
@@
 $Out=M_{w}D$ (3)
-Figure 2: $D$ is obtained by performing edge detection convolution on $T$ and unfolding the output.
+Figure 2: $D^{n}$ is obtained by performing edge detection convolution with dilation ratio of n on $T$ and unfolding the output.
 When $M_{w}$ is influenced by the input $T$, the CDC could exhibit enhanced robustness.
@@
 3.
-Figure 3: The various heads in DATrans employ the edge convolution to get $D_{\sum}$. Each head’s key and value are derived from $D_{\sum}$, while the query is generated from the flattened input.
+Figure 3: The overall structure of DATrans.
 In varied heads of the DATrans, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
@@
 2 to $T$, utilizing varying dilation ratios.
-The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$.
+The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ where $n$ is the dilation ratio.
 Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$.
@@
 $Q^{n}$ $=W^{n}_{Q}O,K^{n}=W^{n}_{K}D^{n},V=W^{n}_{V}D^{n}$ (4)
-where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions, $n$ denotes dilation ratios, and $m$ denotes the number of heads.
+where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions and $m$ denotes the number of heads.
 We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$.
@@
 ALCNet[14] 0.37 92.45 96.08 98.94 2.62 65.10 78.88 93.26 7.90
-AGPCNet[15] 13.36 85.38 92.11 97.88 5.95 67.54 80.64 94.27 4.76
+AGPCNet[15] 13.36 86.43 92.72 98.20 3.70 67.54 80.64 94.27 4.76
 DNANet[7] 4.69 93.73 96.70 99.25 4.55 65.79 79.23 93.60 5.96

### v4

@@
 However, ISTD encounters several challenges, including the tendency for small and dim targets to be obscured by complex backgrounds.
-To address this issue, we propose the Dynamic Attention Transformer Network (DATransNet), which aims to extract and preserve edge information of small targets.
-DATransNet employs the Dynamic Attention Transformer (DATrans), simulating central difference convolutions (CDC) to extract and integrate gradient features with deeper features.
-Furthermore, we propose a global feature extraction module (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the background information.
-We compare the network with state-of-the-art (SOTA) approaches, and the results demonstrate that our method performs effectively.
+To address this issue, we propose the Dynamic Attention Transformer Network (DATransNet), which aims to extract and preserve detailed information vital for small targets.
+DATransNet employs the Dynamic Attention Transformer (DATrans), simulating central difference convolutions (CDC) to extract gradient features.
+Furthermore, we propose a global feature extraction module (GFEM) that offers a comprehensive perspective to prevent the network from focusing solely on details while neglecting the global information.
+We compare the network with state-of-the-art (SOTA) approaches and demonstrate that our method performs effectively.
 Our source code is available at https://github.com/greekinRoma/DATransNet.
@@
 I Introduction
-INFRARED small target detection (ISTD) is critically important for numerous space-based computer vision applications.
-Although ISTD has been extensively studied, it remains a challenge.
-The considerable imaging distances associated with space-based infrared cameras often result in small targets that lack detailed texture and shape information.
-Currently, the ISTD methods can be categorized into model-driven and data-driven methods.
+INFRARED small target detection (ISTD) is vital for various fields. Currently, ISTD methods can be categorized into model-driven and data-driven methods.
 Model-driven methods include three main approaches.
-1) Filter-based methods, such as the Tophat [1], enhance target visibility in complex backgrounds while suppressing background noise.
-2) Methods based on the human visual system (HVS), such as the local contrast measure (LCM)[2] and multi-patch contrast measure (MPCM)[3], leverage human visual perception to improve local information.
-3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch image (IPI)[4], and Rewight infrared patch tensor (RIPT) [5], decompose small targets and backgrounds using low-rank matrices.
-However, these techniques require manual parameter settings and often struggle to adapt to complex scenarios.
-With the advancement of deep learning, data-driven approaches have achieved substantial progress in infrared small target detection (ISTD).
-For example, the Asymmetric Context Modulation (ACM) network [6] introduces an asymmetric feature fusion technique, an alternative to conventional skip connections in U-Net.
-Building on this, DNANet [7] implements a multilayer nested architecture that supports progressive and adaptive interactions among feature layers.
+1) Filter-based methods, such as the Tophat [1].
+2) Methods based on the human visual system (HVS), such as the local contrast measure (LCM)[2] and multi-patch contrast measure (MPCM)[3].
+3) Low-rank matrix decomposition and reconstruction methods, such as infrared patch image (IPI)[4], and reweighted infrared patch tensor (RIPT) [5].
+With the advancement of deep learning, data-driven approaches have achieved substantial progress in ISTD.
+For example, the Asymmetric Context Modulation (ACM) network [6] introduces asymmetric feature fusion, an alternative to conventional skip connections in U-Net.
+The dense nested attention network (DNANet) [7] implements a multilayer nested architecture that supports progressive and adaptive interactions between feature layers.
 Moreover, UIUNet [8] enhances the detection of local target contrasts by integrating multiple U-Net structures and using interactive cross-attention mechanisms for feature fusion.
-Additionally, GSTUnet [9] merges vision Transformer technology with CNNs in the encoder to efficiently extract edge features from small targets.
+Additionally, Gated-shaped TransUnet (GSTUnet) [9] merges Vision Transformer with CNNs in the encoder to learn both global and local information.
+Receptive-field and Direction-induced Attention Network (RDIAN) [10] utilizes different receptive fields and multi-direction-guided attention to enhance the features of targets.
+Attentional Local Contrast Network (ALCNet) [11] designs a bottom-up attention modulation to strengthen the small target characteristics.
+Attention-guided Pyramid Context Network (AGPCNet) [12] divides the image into several patches and computes both global and local associations.
+Yuan et al.
+[13] propose the Spatial-channel Cross Transformer Network (SCTransNet), which uses the Spatial-channel Cross Transformer Blocks (SCTBs) to improve the capacity of global information modelling.
 Generally, these data-driven methods surpass traditional model-driven approaches.
-Some challenges faced by data-driven methods, such as the substantial number of parameters and the potential loss of local feature information.
-Our motivation focus on enhancing the feature information from small targets.
-Although Data-driven methods achieve good performance, they suffer from a substantial number of parameters and potential loss of local information.
-Therefore, we propose some methods to improve the ISTD network, and the new network is called Dynamic Attention Transformer Network (DATransNet).
-DATransNet incorporates two key innovation components to improve feature extraction.
-First, we propose the Dynamic Attention Transformer (DATrans), which utilizes Transformers to simulate the central difference convolution (CDC) and enable the network to retain the detailed information of the inputs in deep layers and give dynamic weights to differences result from differenct directions.
-Second, the global feature extraction module (GFEM) solves the problem of lacking global background perception, thereby improving the ability to obtain contextual information.
+Although data-driven methods achieve excellent performance, they often struggle with limited ability to capture details and weak global perceptual capabilities.
+To address these limitations, we propose a novel framework called the Dynamic Attention Transformer Network (DATransNet).
+DATransNet incorporates two key modules.
+We first introduce the Dynamic Attention Transformer (DATrans), which is good at capturing details.
+Then, the Global Feature Extractor Module (GFEM) gives our model a global perspective over the whole image.
 The main contributions of this letter are as follows.
-Figure 1: The overall structure of DATransNet. The red represents the downsampling module and the blue corresponds to the upsampling module.
+Figure 1: The overall structure of DATransNet. The red represents the upsampling stage (UpStage), and the blue corresponds to the downsampling stage (Stage).
   1. 1.
-    We propose the DATransNet, which utilizes the DATrans to simulate the CDC with dynamic weights, extracting the edge and point information.
+    We propose DATransNet, which utilizes DATrans to extract detailed information by simulating the central difference convolution (CDC) with dynamic weights.
   2. 2.
-    We propose the GFEM to improve the CNN from only focusing on local details to integrating these details with global information.
+    We introduce GFEM to incorporate global information into our network.
   3. 3.
-    We conduct comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that DATransNet outperforms existing methods.
+    We conducted comparative experiments on the IRSTD-1K and NUDT-SIRST datasets, and the results show that DATransNet outperforms many existing methods.
+Figure 2: The first column is the original image, and the following 8 columns are the results of the original image using different edge convolutions across various directions.
 II Methods
-The overall architecture of DATransNet is illustrated in Fig. 1. DATrans is based on U-Net and implements DATrans and GFEM to extract detailed information and contextual features.
+The overall architecture of DATransNet is illustrated in Fig. 1. DATransNet is based on U-Net and implements DATrans and GFEM to extract detailed information and contextual features, respectively.
 II-A Dynamic Attention Transformer (DATrans)
+As shown in Fig.
+2, the targets in ISTD are small and dim and do not have complex texture information.
+Therefore, the details in the image are vital for ISTD.
+An effective approach to extracting the details is to take advantage of the differences between local pixels and their surroundings, such as CDCs.
 CDCs are defined in Eq.
-1 and are employed to extract edge and gradient features.
+1.
+$out_{\hat{c}xy}=\sum^{c_{inp}-1}_{i=0}\sum^{7}_{j=0}w_{\hat{c}ijxy}(b_{ijxy}-o_{ixy})\$ (1)
 Here, $c_{inp}$ represents the number of input channels, $o_{ixy}$ denotes the input value at the $i$-th channel and position $(x,y)$, and $b_{ijxy}$ refers to the $j$-th surrounding value around $o_{ixy}$.
 $out_{\hat{c}xy}$ denotes the output of the $\hat{c}$-th channel at position $(x,y)$.
-$w_{\hat{c}ijxy}$ is the weight applied to the difference $(o_{ixy}-b_{ijxy})$, which contributes to determining $out_{\hat{c}xy}$.
-A significant limitation of CDCs is the use of fixed weights for different directions, meaning the importance of each direction remains constant across different images.
-$out_{\hat{c}xy}=\sum^{c_{inp}-1}_{i=0}\sum^{7}_{j=0}w_{\hat{c}ijxy}(b_{ijxy}-o_{ixy})\$ (1)
-We can transform Eq.
-1 into a fusion of Eq.
+$w_{\hat{c}ijxy}$ is the weight applied to the difference $b_{ijxy}-o_{ixy}$.
+The importance of different edge-convolution results varies from image to image.
+As shown in Fig.
+2, the target in the upper image of the last column is clearly distinct, while that in the bottom row appears blurry.
+So, the importance of difference results in various directions is varied for different images.
+The weights $w_{\hat{c}ijxy}$ of the CDCs should be dynamically adjusted in response to changes in the input images.
+We divide Eq. 1 into a fusion of Eq. 2, Eq. 3, and Eq. 4.
+Eq.
+2 and Eq.
+3 describe the process of extracting difference features $T^{n}\in\mathbb{R}^{8c_{inp}\times(wh)}$, as shown in Fig.
+3.
+We apply edge convolutions with a dilation ratio of $n$ to get $D^{n}\in\mathbb{R}^{8c_{inp}\times w\times h}$ from the input image $I\in\mathbb{R}^{c_{inp}\times w\times h}$, as outlined in Eq.
+2.
+The $Diff$ refers to performing eight edge convolutions on the input image.
+In Eq.
+3, the $Flatten$ reshapes $D^{n}$ into $T^{n}$.
+In addition, $w$ represents the input width, $h$ denotes the input height, and $c_{inp}$ denotes the number of input channels.
+In the process, we convert the spatial differences between eight neighborhoods to the dimension of the channel.
+Then we use the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{inp}}$ to obtain the output $Out\in\mathbb{R}^{8c_{inp}\times(wh)}$, as shown in Eq.
+4, where $Reshape$ refers to the transformation process transforming $M_{w}T^{n}$ into $Out^{n}\in R^{c_{o}\times w\times h}$ and $c_{o}$ is the number of output channels.
+$D^{n}=Diff(I)$ (2)
+$T^{n}=Flatten(D^{n})$ (3)
+$Out^{n}=Reshape(M_{w}T^{n})$ (4)
+Figure 3: $T^{n}$ is derived by applying edge detection convolutions with a dilation ratio of $n$ to the image $I$, followed by flattening the output.
+When $M_{w}$ in Eq.
+4 changes with the input image $I$, the CDCs may also vary accordingly.
+Based on the analysis, we introduce the DATrans to simulate CDCs with dynamic weights for different directions.
+The structure of DATrans is shown in Fig.
+4.
+To improve the detection capability for a variety of targets, we utilize varying dilation ratios across different heads in DATrans.
+For each head, the difference feature ($T^{n}$) extraction process follows the same approach as that described in Eq.
 2 and Eq.
 3.
-The process described in Eq.
-2 is illustrated in Fig.
-2.
-As shown in Fig.
-2, we use edge convolution to extract the difference values from the input $T\in\mathbb{R}^{c_{inp}\times w\times h}$ and reshape them to produce a local difference matrix $D\in\mathbb{R}^{8c_{inp}\times wh}$.
-In Eq.
-3, we utilize the weight matrix $M_{w}\in\mathbb{R}^{c_{o}\times 8c_{inp}}$ and the local difference matrix $D$ to obtain the final result $Out$.
-Here, $w$ denotes the width of the input, $h$ represents the input height, and $c_{o}$ signifies the number of output channels.
-$D=reshape(Conv_{3\times 3}(T))\$ (2)
-$Out=M_{w}D$ (3)
-Figure 2: $D^{n}$ is obtained by performing edge detection convolution with dilation ratio of n on $T$ and unfolding the output.
-When $M_{w}$ is influenced by the input $T$, the CDC could exhibit enhanced robustness.
-Based on the analysis above, we introduce the DATrans to simulate CDC and extract the details using dynamic weights from low-level layers for ISTD.
-The structure of DATrans is shown in Fig.
-3.
+We treat each channel with 8 surrounding information as a token.
 Figure 4: The overall structure of DATrans.
-In varied heads of the DATrans, we employ different unfolding methods based on edge detection convolutions in Fig. 2 with varying dilation and reshape operations to get the unfolding .
-In the various heads of DATrans, we apply different edge detection convolutions from Fig.
-2 to $T$, utilizing varying dilation ratios.
-The outputs from these convolutions are then reshaped into $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ where $n$ is the dilation ratio.
-Additionally, the input $T$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$.
-The token $O$ serves as the query, and the local difference matrix $D^{n}\in\mathbb{R}^{8c_{inp}\times wh}$ functions as both the key and the value.
-$Q^{n}$ $=W^{n}_{Q}O,K^{n}=W^{n}_{K}D^{n},V=W^{n}_{V}D^{n}$ (4)
-where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the reshaped weights of convolutions and $m$ denotes the number of heads.
-We utilize the $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$.
-$V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ denotes the value of the transformer.
+Furthermore, the input $I$ is reshaped to form $O\in\mathbb{R}^{c_{inp}\times wh}$, as shown in Eq.
+5.
+$O$ serves as the query token.
+The key and value tokens come from $T^{n}\in\mathbb{R}^{8c_{inp}\times wh}$.
+$O=Flatten(I)$ (5)
+$Q^{n}$ $=W^{n}_{Q}O,K^{n}=W^{n}_{K}T^{n},V=W^{n}_{V}T^{n}$ (6)
+where $W^{n}_{Q}\in\mathbb{R}^{\frac{c_{o}}{m}\times c_{inp}}$, $W^{n}_{K}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$, $W^{n}_{V}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times 8c_{inp}}$ are the learnable weights and $m$ denotes the number of heads.
+We use $Q^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ and $K^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ to produce the attention matrix $M^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times\frac{8c_{inp}}{m}}$.
+$V^{n}\in\mathbb{R}^{\frac{8c_{inp}}{m}\times wh}$ is multiplied by $M^{n}$.
 The result $Out^{n}\in\mathbb{R}^{\frac{c_{o}}{m}\times wh}$ is defined as follows:
@@
 $=M^{n}V^{n}=(M^{n}W_{V}^{n})D^{n}=M^{n}_{mix}D^{n}$
-where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{inp}}$ denotes the dynamic matrix derived from a attention matrix $M^{n}$ and a fixed convolution weight matrix $W_{V}^{n}$.
+where $M^{n}_{mix}\in\mathbb{R}^{\frac{c_{o}}{m}\times 8c_{inp}}$ denotes the dynamic matrix derived from an attention matrix $M^{n}$ and a learnable weight matrix $W_{V}^{n}$.
+$Norm$ refers to Instance Normalization, which normalizes the similarity matrix for each instance on the similarity maps, ensuring smooth gradient propagation.
 According to Eq.
-3, $Out^{n}$ is the equivalent result of the CDC with dynamic weights and a dilation ratio of n.
-$Norm$ is the Instance Normalization.
-At last, we concatenate the results in varied heads, reshape them and use $1\times 1$ convolution to fuse them as shown in the Equation below:
+4 and 7, $Out^{n}$ is the equivalent result of the CDC with dynamic weights.
+At last, we concatenate the results of varied heads which have different dilation ratios, reshape them and use a $1\times 1$ convolution to fuse them to get the final output $Out$.
+In addition, $m$ is the number of heads.
 $Out=Conv_{1\times 1}(Reshape(Out^{1},...,Out^{m}))$ (8)
-Overall, the gradient information from features at different scales can be fully extracted, enhancing the model’s ability to adapt to complex backgrounds.
 II-B Global Feature Extraction Module (GFEM)
-Background information is crucial alongside detailed features in ISTD.
-However, perceiving background relies on global features, which the DATrans struggles to handle effectively.
-To address this limitation and enhance global information perception, we propose the GFEM, as illustrated in Fig.
-4.
-In GFEM, we utilize the attention mechanism to facilitate the transmission of information across distant pixels.
-This module includes two key aspects to broaden perception: First, we employ the non-local attention mechanism [10] to compute the spatial attention, enabling connectivity between pixels over long distances.
-Second, we incorporate a squeeze-and-excitation block [11] to enhance channel-wise perception.
-Finally, we concatenate the outputs from the global spatial and channel attention modules and apply convolution to fuse them, enabling GFEM to capture comprehensive feature representations.
-Subsequently, each feature map acquires both spatial and channel global receptive fields, which are vital for effective object detection.
+Background information is crucial alongside the detailed features of small targets in ISTD.
+However, background information is based on a global perspective of the whole image.
+So, we propose the GFEM, as illustrated in Fig.
+5.
 Figure 5: The structure of the GFEM using the Non-local (Non-local Attention Module) to capture global spatial features and SE Block (Squeeze-and-Excitation Block) to extract global channel information.
+In GFEM, we incorporate the attention mechanism to provide our model with a broader perspective.
+This module includes three key steps: First, we utilize the non-local attention mechanism [14] to compute spatial attention over the deepest feature map, which is significantly smaller in size compared to the input image.
+Initially, we apply a $1\times 1$ convolution to reduce the number of channels from $c$ in $X\in\mathbb{R}^{c\times w\times h}$ to $c^{\prime}$ in $Q\in\mathbb{R}^{c^{\prime}\times w\times h}$, $K\in\mathbb{R}^{c^{\prime}\times w\times h}$, and $V\in\mathbb{R}^{c^{\prime}\times w\times h}$.
+Besides, $w$ and $h$ denote the width and height of the feature map input to the non-local module.
+Then, we reshape the feature maps and compute $Y\in\mathbb{R}^{c\times w\times h}$ according to the following equation:
+$Y=Conv_{1\times 1}(Reshape(V^{\prime}\cdot Softmax(Q^{\prime}\cdot(K^{\prime})^{T})))$
+where $Q^{\prime}\in\mathbb{R}^{c^{\prime}\times(wh)}$, $K^{\prime}\in\mathbb{R}^{c^{\prime}\times(wh)}$, and $V^{\prime}\in\mathbb{R}^{c^{\prime}\times(wh)}$ are the results of reshaping $Q$, $K$, and $V$.
+$Softmax$ is softmax layer, and $Conv_{1\times 1}$ is a convolution with a $1\times 1$ kernel, to increase the number of channels.
+Second, we incorporate a squeeze and excitation block [15] to enhance channel-wise perception.
+Finally, we concatenate the results from the global spatial and channel attention modules and employ convolution to fuse them, enabling GFEM to capture comprehensive feature representations.
+Subsequently, the output feature map acquires both spatial and channel global receptive fields, which are vital for effective object detection.
 II-C Loss Function
-The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, where the predictions of DATransNet are mapped into the range of 0 and 1, using the Sigmoid function as shown in Eq.
-7.
-$p_{i,j}=\frac{1}{1+e^{-v_{i,j}}}$ (7)
-where $p_{i,j}$ denotes the Sigmoid output of $v_{i,j}$, the model’s prediction at pixel $(i,j)$. Moreover, the SoftIoU loss is defined as:
+The loss function utilized in our model training is the soft intersection over union (SoftIoU) loss, as shown in Eq. 9.
 $loss=1-\frac{\sum_{i,j}p_{i,j}\cdot g_{i,j}}{\sum_{i,j}p_{i,j}+\sum_{i,j}g_{i,j}-\sum_{i,j}p_{i,j}\cdot g_{i,j}}$ (9)
-In this equation, $g_{i,j}$ denotes the ground truth value at pixel $(i,j)$, typically 0 or 1.
-Figure 5: Visual results from varied data-driven methods.
-Red, green, and yellow boxes label detected targets, missed targets, and false alarms.
-The Red bounding box are zoom in, facilitating to auxiliary our to evaluation detection results.
-Our DATransNet can achieve high probability detection rate, and achieve precision localization and segmentation of their shapes.
+where $g_{i,j}$ and $p_{i,j}$ denote the ground truth and the output of our network at the $(i,j)$, respectively.
+Figure 6: Visual results from varied data-driven methods. The red, green and yellow boxes represent detected targets, missed targets, and false alarms.
 III Experiment and Analysis
-This section first describes the experimental details and the evaluation metrics.
-Following with a series of ablation studies to verify the proposed modules.
-Finally, a comprehensive comparative experiments with other methods on qualitative and quantitative results demonstrate that our approach outperforms existing state-of-the-art (SOTA) methods.
+This section describes the experimental details and the evaluation metrics, followed by a series of ablation studies to verify the proposed modules.
+Finally, a comprehensive comparative experiment with other methods on qualitative and quantitative results demonstrates that our approach outperforms existing state-of-the-art (SOTA) methods.
 III-A Dataset and Evaluation Metrics
 The datasets used for evaluating the DATransNet are NUDT-SIRST[7] and IRSTD-1K[16].
-The evaluation metrics include mean intersection over union ($mIoU$), F-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$).
-Meanwhile, we plot the ROC curves of various methods to evaluate the robust of DATransNet.
+The NUDT-SIRST dataset consists of 1327 images whose resolution is 256×256, with a split of 332 images for testing, 332 for validation, and 663 for training.
+The IRSTD-1K dataset contains 1001 images with a resolution of 512×512.
+We select 101 images for testing, 100 images for validation, and 800 images for training.
+The evaluation metrics include mean intersection over union ($mIoU$), F1-measure ($F_{1}$), the probability of detection ($P_{d}$), and false alarm rate ($F_{a}$).
 III-B Experimental Environment and Parameter Settings
-Our model is implemented by PyTorch framework, then training on NVIDIA GeForce RTX 4080 GPUs.
+All models are built on the PyTorch framework and are trained on an NVIDIA GeForce RTX 4080 GPU.
 We employ the Adam optimizer, starting with an initial learning rate of $5\times 10^{-4}$.
-This learning rate is reduced to $5\times 10^{-5}$ at epoch 200 and further decrease to $5\times 10^{-6}$ We use a batch size of 4 for training and run the model for 400 epochs.
+This learning rate is reduced to $5\times 10^{-5}$ at epoch 200 and further decreases to $5\times 10^{-6}$.
+The batch size is 4, and there are 400 epochs.
 III-C Ablation Study
-III-C1 Module-wise Performance Gain
-In this ablation study, we begin with the most primitive Baseline, Which includes only feature extraction and upsampling with in the U-Net framework.
-Then, We test the performance of the DATransNet w/o each module, the results are shown in Tab.
+III-C1 Studies of Module-wise Performance Gain
+In this ablation study, we begin with the baseline.
+Then, we test the performance of the DATransNet’s module, and the results are shown in Tab.
 I.
-The results illustrate the effectiveness of each module.
-The performance of our network improves progressively with the inclusion of DATrans and GFEM.
-DATrans, which focuses on local difference features, enhances the model’s performance more effectively than GFEM, which primarily relies on global perception.
-In ISTD, the capability to learn gradient information is more critical than access to global context.
+The integration of DATrans and GFEM leads to a progressive enhancement in the performance of our network..
 TABLE I: Studies Of Different Components on NUDT-SIRST (The best results are bold)
@@
 Tab.
-II demonstrates the efficacy of this strategy, as models with diverse dilation ratios outperform those with a uniform rate.
-Furthermore, networks with dilation ratios of 1 and 3 yield superior results compared to those with other dilation ratios.
-Therefore, the dilation ratio significantly impacts target detection, and we choose the ratios of 1 and 3 as the dilation ratios for our network.
+II demonstrates the efficacy of this strategy, as models with diverse dilation ratios outperform those with a single rate.
+Furthermore, networks with dilation ratios of 1 and 3 yield superior results.
 TABLE II: Studies on different dilation rates on NUDT-SIRST (The best results are bold)
@@
 TABLE III: Studies on GFEM on NUDT-SIRST (The best results are bold)
-Non-local SE Block mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
-$\checkmark$ -- 94.69 97.27 98.73 2.57
--- $\checkmark$ 94.53 97.19 98.23 1.75
-$\checkmark$ $\checkmark$ 94.93 97.39 99.04 2.00
+Non-local SE Block Params(M) GFLOPs(G) mIoU(%) $F_{1}(\%)$
+-- -- 3.70 10.82 94.25 97.03
+$\checkmark$ -- 4.03 10.89 94.69 97.27
+-- $\checkmark$ 4.02 10.89 94.53 97.19
+$\checkmark$ $\checkmark$ 4.04 10.90 94.93 97.39
+TABLE IV: Quantitative Comparsion With Different Methods on NUDT-SIRST and IRSTD-1K (The best results are bold, second best results are underline.)
+Model Metrics NUDT-SIRST IRSTD-1K
+Params(M) mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$ mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $F_{a}(10^{-6})$
+ACM[6] 0.40 70.97 82.99 97.67 7.33 64.09 78.11 88.55 17.21
+RDIAN[10] 0.90 87.78 93.47 97.88 9.67 65.68 79.30 91.24 10.31
+ALCNet[11] 0.37 92.45 96.08 98.94 2.62 65.05 78.60 90.57 19.42
+DNANet[7] 4.69 93.73 96.70 99.25 4.55 66.73 80.08 89.22 7.97
+UIU-Net[8] 50.54 94.11 96.96 97.98 0.74 67.94 80.90 90.57 24.44
+SCTransNet[13] 11.19 93.83 96.77 98.20 1.56 65.83 79.40 90.57 14.50
+Ours 4.04 94.93 97.39 99.04 2.00 68.56 81.34 93.60 24.96
 III-C3 Studies of GFEM
-As mentioned in III, we conduct the experiments on the modules in the GFEM. As we can see, the network combining the two modules performs better.
+As mentioned in III, we conduct the experiments on the modules in the GFEM. The network combining the two modules performs better with a slight increase in computational complexity.
 III-D Comparsion with State-of-the-art (SOTA) Methods
-To evaluate the effectiveness of our proposed model, we compare it with several SOTA methods on NUDT-SIRST and IRSTD-1K datasets, including SCTransNet [16], UIUNet [8], ACM [6], ALCNet [14], AGPCNet [15], RDIAN [13], and DNANet [7].
-As shown in Tab.
-IV, and our method performs excellently.
-We utilize DATransNet with dilation ratios of 1 and 5 as the baseline models for comparative evaluation.
-The comparison experiments include qualitative results and quantitative results, receptively.
-III-D1 Qualitative Result
-As shown in Fig.
-5, our model preserves the shape of targets more closely to ground truth with low missed detections (green boxes) and false alarms (yellow boxes) compared to other methods, achieve satisfied results.
-TABLE IV: Quantitative Comparsion With Different Methods on NUDT-SIRST and IRSTD-1K (The best results are bold, second best results are underline.)
-Module Metrics NUDT-SIRST IRSTD-1K
-Params(M) mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$ mIoU(%) $F_{1}(\%)$ $P_{d}(\%)$ $P_{a}(10^{-6})$
-ACM[6] 0.40 70.97 82.99 97.67 7.33 63.78 77.91 93.93 5.81
-RDIAN[13] 0.90 87.78 93.47 97.88 9.67 64.08 78.12 93.34 5.30
-ALCNet[14] 0.37 92.45 96.08 98.94 2.62 65.10 78.88 93.26 7.90
-AGPCNet[15] 13.36 86.43 92.72 98.20 3.70 67.54 80.64 94.27 4.76
-DNANet[7] 4.69 93.73 96.70 99.25 4.55 65.79 79.23 93.60 5.96
-UIU-Net[8] 50.54 94.11 96.96 97.98 0.74 68.06 80.91 92.59 5.85
-SCTransNet[16] 11.19 93.83 96.77 98.20 1.56 67.97 80.90 93.26 6.24
-Ours 4.46 94.93 97.39 99.04 2.00 68.26 81.11 95.95 4.42
+We compare it with several SOTA methods on NUDT-SIRST and IRSTD-1K, including SCTransNet [13], UIUNet [8], ACM [6], ALCNet [11], AGPCNet [12], RDIAN [10], and DNANet [7].
+In addition, deep supervision is not employed to ensure equality in any of the networks.
+We used DATransNet with dilation ratios of 1 and 3 for comparative evaluation.
+The results of the experiments include qualitative and quantitative results.
+III-D1 Qualitative Results
+As shown in Fig. 6, our model preserves the shape of targets more closely to ground truth with low missed detections and false alarms.
 III-D2 Quantitative Results
 As shown in the Tab.
-IV, our method achieves impressive performance, including an $mIoU$ of $94.93\%$, an $F1$ score of $97.39\%$, and a $Pd$ of $99.04\%$, a $Pa$ of $2.00\times 10^{-6}$.
+IV, our method achieves impressive performance, including a $mIoU$ of $94.93\%$, an $F1$ score of $97.39\%$, and a $Pd$ of $99.04\%$, a $F_{a}$ of $2.00\times 10^{-6}$.
 Similarly, on the IRSTD-1K dataset, our method is a leading network.
-Figure 6: ROC curves of varied networks on NUDT-SIRST datasets.
-Furthermore, we used the ROC curve to analyze the performance of various models, as depicted in Fig.
-6.
-Although UIU-Net achieves commendable results in $mIoU$, its AUC scores are comparatively weak.
-In contrast, ALCNet performance shows a significant divergence, with subpar outcomes in $mIoU$ but relatively strong performance in ROC metrics.
-The ROC curve corroborates the findings in Tab.
-IV.
-The model proposed in this article consistently shows good performance in both ROC and $mIoU$.
+Figure 7: ROC curves of varied networks on NUDT-SIRST and IRSTD-1K datasets.
+Furthermore, we used the ROC curve to analyze the performance of various models, as depicted in Fig. 7. The model proposed in this article consistently shows good performance in both ROC and $mIoU$.
 IV Conclusion
 In this Letter, we propose an ISTD network that enhances detection performance.
-The DATrans was proposed to enhance the local gradient feature extraction and detection performance of the network.
-Additionally, we propose the GFEM, which focuses on global perception across the entire feature map and learns relationships between distant pixels, preventing DATrans from overlooking background information due to its emphasis on local gradient differences.
-Then the local and global feature are fused.
-We conducted extensive experiments indicating satisfied results with fewer parameters.
-The model is a back-box approach because its capability for dynamic weight adjustments lacks mathematical proof.
+DATrans was proposed to enhance the network’s local gradient feature extraction and detection performance.
+Additionally, we propose the GFEM, which focuses on global perception across the entire feature map and learns relationships between distant pixels.
+We have conducted extensive experiments, indicating satisfied results with fewer parameters.
+The model is still a black-box approach because its capability for dynamic weight adjustments lacks mathematical proof.
 Moreover, it suffers from false alarms when processing high-noise images.
-To address these limitations, we aim to improve the model in two aspects.
+To address these limitations, we aim to improve the model in two ways.
 At first, we plan to improve the interpretability of the network.